A hot lump of 38.4 g of aluminum at an initial temperature of 71.6 °C is placed in 50.0 mL H2O initially at 25.0 °C and allowed to reach thermal equilibrium. What is the final temperature (C) of the aluminum and water, given that the specific heat of aluminum is 0.903 J/(g·°C)? Assume no heat is lost to surroundings
Answer #1

mass of H2O   = volume * density

                       = 50*1 = 50g

heat lose of Al                  =              heat gain of H2O

mcT                               =              mcT

38.4*0.903*(71.6-t)           =              50*4.184*(t-25)

t   = 31.6250C

The final temperature is 31.6250C